10 July, 2001
Gray, cold, windy day. Temperature is well below the normal 45 degrees. It is currently below freezing with the wind chill taken into account. We ventured out to the site this afternoon to do tower transects. I was working on the plant identification and Lety and Michelle were taking soil moisture and thaw depth measurements. These are done on a weekly basis because they have a direct effect on the carbon exchange system.
With the wind blowing we managed to stay out for a few hours and completed enough to keep us busy in the lab. I have started recording the 1998 data onto an excel sheet and will get it into a GIS database with Joe's help.
Other than that it was a slow but steady day today. Hyojung and Spring were in the lab all day as was Glen crunching numbers. Rommel and Joe are tuning up the plane today. Rommel didn't fly today because of the foggy weather.
Contact the TEA in the field at .
If you cannot connect through your browser, copy the TEA's e-mail address in the "To:" line of your favorite e-mail package.